OpenStudy (anonymous):

what is the inverse of this function i(x)=3-4/2x+1 Please help Can anybody help me please I am stressed to the max LOL

OpenStudy (anonymous):

\[i(x)=3 -\frac{4}{2y+1}\]is the same as saying \[x=3-\frac{4}{2y+1}\] What you're trying to find in an inverse is basically switching the positions of the x and y values and making the x-value the subject again. So... \[y=3-\frac{4}{2x+1}\] Now we make x the subject. Try and do that and someone else may get back to you as I have to leave. Sorry for the inconvenience.

OpenStudy (anonymous):

the last equation you have, is that what you want the inverse of because that is the same as what you have in the original question.

OpenStudy (anonymous):

i want the inverse of i(x)=\[3-4\div2x+1\]

OpenStudy (anonymous):

ok, so in that equation, switch the x and y. so you get \[x=3-\frac{ 4 }{2y+1}\]

OpenStudy (anonymous):

then solve for y. so what is the first thing you will do?

OpenStudy (anonymous):

x-3

OpenStudy (anonymous):

correct. so then you have x-3=4/(2y+1). Now what you want to do is get y out of the bottom

OpenStudy (anonymous):

isn't the rhs negative? Do I then divide by y?

OpenStudy (anonymous):

yes it is, sorry i thought i put the sign in there...you want to multiply both sides by (2y+1) so that you get (2y+1)(x-3)=4

OpenStudy (anonymous):

negative 4, sorry again

OpenStudy (anonymous):

now that you have (2y+1)(x-3)=-4 how do you get to y=?

OpenStudy (anonymous):

do i expand the brackets?

OpenStudy (anonymous):

no, the first step is to get (x-3) to the rhs

OpenStudy (anonymous):

so divide by x-3

OpenStudy (anonymous):

yes, and then all you have left is 2y+1=-4/(x-3)

OpenStudy (anonymous):

so then -1 and get 2y = (-4/(x-3)) - 1 ?

OpenStudy (anonymous):

yes, and then divide everything by 2

OpenStudy (anonymous):

\[y= -2/((x-3)/2)- 1/2\]

OpenStudy (anonymous):

I think that is right, just to be sure, I got \[y=\frac{ -2 }{ x-3 }+1/2\]

OpenStudy (anonymous):

is this for written out hw or online hw?

OpenStudy (anonymous):

so the denominator is not divided by 2 as well?

OpenStudy (anonymous):

written out :)

OpenStudy (anonymous):

no because it is a fraction divided by a number, it is really \[(\frac{ 1 }{ 2 })(\frac{ -4 }{ x-3 })\]

OpenStudy (anonymous):

ok, and have you ever heard of wolframalpha.com?

OpenStudy (anonymous):

yes but I needed step by step to help me understand the process

OpenStudy (anonymous):

ok, you might want to check it against what my answer was just to be sure, and if there is a different answer, let me know and ill figure it out

OpenStudy (anonymous):

No I am happy with this as my asymptotes now match :) Thank you so very very much xoxo

OpenStudy (anonymous):

ok, no problem. Hope you do well on your hw!!

OpenStudy (anonymous):

Thanks again :)